# FarsInstruct: Empowering Large Language Models for Persian Instruction Understanding

### v3

@@ Abstract
Instruction-tuned large language models have demonstrated remarkable capabilities in following human instructions across various domains. However, their proficiency remains notably deficient in many low-resource languages. To address this challenge, we begin by introducing FarsInstruct: a comprehensive instruction dataset designed to enhance the instruction-following ability of large language models specifically for the Persian language—a significant yet underrepresented language globally. FarsInstruct encompasses a wide range of task types and datasets, each containing a mix of straightforward to complex manual written instructions, as well as translations from the Public Pool of Prompts, ensuring a rich linguistic and cultural representation. Furthermore, we introduce Co-CoLA, a framework designed to enhance the multi-task adaptability of LoRA-tuned models. Through extensive experimental analyses, our study showcases the effectiveness of the FarsInstruct dataset coupled with training by the Co-CoLA framework, in improving the performance of large language models within the Persian context. As of the current writing, FarsInstruct comprises 197 templates across 21 distinct datasets, and we intend to update it consistently, thus augmenting its applicability. [^1] https://huggingface.co/datasets/PNLPhub/FarsInstruct
@@ 1 Introduction
The modern era of artificial intelligence is marked by numerous breakthroughs, among which is the rise of large language models (LLMs), such as GPT4 (openai2024gpt4), Llama3 (dubey2024llama3herdmodels) and PaLM (chowdhery2022palm). Instruction-tuning emerges as a vital technique in the evolution of language models, involving training a model on a wide range of tasks described through natural language instructions. This method diverges from traditional task-specific fine-tuning and adapts the model’s behavior to respond to user queries with relevant and helpful answers. This technique offers a more generalized and versatile approach to model training, thus contributing significantly to the advancement of LLMs.
@@
Despite the steady progress of instruction-tuned language models, a persistent limitation remains: their difficulty in capturing the nuanced complexities of low-resource languages. This critical challenge stems from the significant gap in the availability of high-quality instruction datasets tailored to these languages. wang2023far highlights this concern, demonstrating that datasets lacking sufficient multilingual diversity can cause models to lose previously learned multilingual capabilities, leading to performance degradation. Moreover, translating English-centric datasets offers only partial solutions due to several inherent limitations (naous2024having; ramesh-etal-2023-fairness; vanmassenhove-etal-2021-machine). While efforts have been made to compile extensive multilingual instruction datasets (wang-etal-2022-super; singh2024aya; muennighoff2022crosslingual), gaps remain in creating diverse and complex prompts for languages like Persian compared to other languages.
@@
In this study, we propose FarsInstruct, a comprehensive human-annotated instruction dataset created from existing Persian NLP datasets. It includes a mixture of manually written instructions ranging from basic to proficient language levels, alongside translations from the Public Pool of Prompts (P3) (sanh2022multitask), which is a collection of prompted English datasets. To ensure the diversity and representativeness of FarsInstruct, we developed 197 prompt templates derived from 21 distinct public datasets. Each prompt template comprises an input template and a target template, both of which function to extract relevant data fields from their respective datasets and reformat them into a unified structure designed for the instruction-tuning objective. For example, in the case of a Textual Entailment dataset containing the fields Premise, Hypothesis, and Label, an input template might be: ”Can the hypothesis be concluded from the premise? Premise: {Premise}, Hypothesis: {Hypothesis}”, while a corresponding target template could be ”The answer is: {label}”.
@@
The collected public datasets encompass ten different task categories: Text Summarization, Textual Entailment, Text Classification, Sentiment Analysis, Word Sense Disambiguation, Query Paraphrasing, Question Answering, Reading Comprehension, Named Entity Recognition (NER), and Translation. Figure 1 depicts an instance of a prompt within our dataset after applying its respective template. A detailed overview of the FarsInsturct dataset is provided in Section LABEL:sec:farsinstruct.
@@
Figure 1: An example of the prompts utilized in the training process. The Persian version of the prompt is employed for training purposes, while the translated English version is provided to enhance comprehension. The instruction component is highlighted in black, the data fields are marked in orange, and the target answer is indicated in gray. In Appendix LABEL:sec:appendix_prompts, this example is shown in the PromptSource environment.
@@
Additionally, parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) (hu2021lora), not only face challenges in multi-task settings but are also prone to catastrophic forgetting (wang2023multilora; li2024mixlora; kalajdzievski2024scaling). To address these issues, we propose Co-CoLA, a novel integration of CoLA (xia2024chain) with rehearsal training (kirkpatrick2017overcoming). More specifically, we adopt an iterative optimization framework that merges learned low-rank matrices into the model parameters and reinitializes optimization for new LoRA modules. At each iteration, we retrain a subset of data from previously learned tasks, mixing it with the current task’s data during training. This periodic revisiting of earlier tasks ensures that the model retains performance across both old and new tasks, all while preserving computational efficiency. Section LABEL:sec:method presents an in-depth explanation of the Co-CoLA method.
@@
In summary, our contributions to advancing Persian instruction understanding are threefold: (1) We present FarsInstruct, a comprehensive human-annotated instruction dataset for Persian, covering varied and representative tasks for different categories such as text summarization, named entity recognition, and translation. (2) We introduce Co-CoLA, a method that combines CoLA with rehearsal training to mitigate catastrophic forgetting in multi-task learning. (3) We release FarsInstruct as an open-source resource, with a commitment to its continued expansion to include a broader range of tasks and modalities.
@@ 2 Related work
Instruction-tuning: Instruction tuning refers to the process of training language models using specific input-output pairs derived from diverse data sources. This approach enhances the ability of a pre-trained LLM to interpret and respond to a wide range of human requests expressed in natural language. Instruction datasets used for this purpose are typically created in one of three ways: (1) manually created by researchers from existing NLP datasets (wang-etal-2022-super; wei2021finetuned), (2) synthesized by prompting proprietary models with a small, seed dataset (alpaca; wang2022self; honovich-etal-2023-unnatural), or (3) generated entirely from scratch, involving human-written prompt-response pairs (DatabricksBlog2023DollyV2; kopf2024openassistant). In this work, we adopt the first approach to develop FarsInstruct. Previous works such as FLAN (wei2021finetuned) and P3 (sanh2022multitask) have been instrumental in advancing instruction dataset creation. FLAN encompasses over 60 NLP datasets, while P3 features more than 2,000 prompts from 177 datasets, each significantly contributing to the field. SuperNaturalInstruction (wang-etal-2022-super) further advanced the field by assembling a comprehensive benchmark featuring 1,616 expert-written NLP tasks, covering 76 unique task types, and extending support to multiple languages. xP3 (muennighoff2022crosslingual) expanded on P3’s groundwork by including content from 46 languages, adding new tasks like Translation and Program Synthesis that P3 had not tackled. Similarly, Aya (singh2024aya) represents a major multilingual effort, featuring an extensive dataset of 513 million instances across 114 languages. This was achieved through a global collaboration involving fluent speakers who contributed instructional content. Our dataset distinguishes itself from these collections in its depth and adaptability, especially with the inclusion of more challenging tasks in Persian, offering a high level of detail not found in many multilingual efforts. While most such projects primarily use machine translations and cover a narrow range of tasks, our dataset presents a wide array of culturally and linguistically rich tasks.
@@
Parameter effecient fine-tuning: Conventional full-parameter fine-tuning becomes computationally impractical as model size and the number of downstream tasks increase. To address this challenge, recent advancements in PEFT methods advocate for training only a small subset of parameters while leaving the majority of pre-trained model parameters intact. One of the most widely utilized paradigms in PEFT is Low-Rank Adaptation (LoRA) (hu2021lora). LoRA modifies only a small, low-rank portion of the model’s weights by incorporating low-rank matrices into the model’s weights during the training process. Despite the significant computational advantage of LoRA, it falls short in multi-task adaptation, Additionally, kalajdzievski2024scaling demonstrated that PEFT techniques, including LoRA, remain vulnerable to catastrophic forgetting, where models lose previously acquired knowledge when fine-tuned on new tasks. MultiLoRA (wang2023multilora) addresses the limitations of LoRA by reducing the dominance of top singular vectors, horizontally scaling LoRA modules, and altering the initialization of adaptation matrices, which leads to improved performance across multiple tasks with minimal additional parameters. MixLoRA (li2024mixlora) introduces multiple LoRA-based experts within a frozen pre-trained model using a top-k routing strategy to efficiently distribute tasks, independently configure attention layer adapters, and apply auxiliary load balance loss, significantly enhancing performance while reducing GPU memory consumption and training latency. Further, CoLA(xia2024chain) introduces an iterative optimization framework designed to improve the fine-tuning of LLMs by employing multiple iterations of LoRA. In this paper, we design Co-CoLA to address the issue of catastrophic forgetting, while ensuring an effective multi-task adaption.
@@ 3 FarsInstruct Dataset
With about 130 million [^2] https://en.wikipedia.org/wiki/Persian_language speakers, Persian — also referred to as Farsi in Iran — is an important language in the Middle East and Central Asia. FarsInstruct represents a project to provide a comprehensive public instruction dataset for the Persian community. As of this writing, FarsInstruct has 197 carefully designed and created prompt templates for 21 already-published public datasets and some translations from existing prompted datasets. Unlike multilingual collections focusing on common tasks such as Text Summarization and Question Answering, FarsInstruct introduces more task types, including Named Entity Recognition and Word Sense Disambiguation. The creation procedure, statistics, task augmentation, and quality of the dataset are covered in detail in the following subsections. Additional illustrations and tables are provided in the Appendix LABEL:dataset_detailsLABEL:sec:appendix_dataset_assetsLABEL:sec:appendix_prompts.
@@ 3.1 Dataset Construction
To efficiently create a large collection of prompts, we primarily utilized PromptSource (bach2022promptsource), an open-source tool designed for creating, sharing, and managing prompts for NLP tasks. A key design choice in bach2022promptsource is the use of Jinja2333https://jinja.palletsprojects.com/en/3.1.x/ as a templating language, providing the flexibility crucial for crafting clear and effective prompts. Each dataset has multiple prompt templates, each of which consists of an input and a target template. These templates map raw data fields into natural language, structuring both the input and target sequences. Practically, templates allow users to mix arbitrary text with data fields. We refer to the text within the input template that guides the model’s behavior as ”Instruction”. Additionally, each prompt template documents essential metadata, including evaluation metrics and the language used.
The PromptSource toolkit offers an interface for interactively writing prompts on datasets. However, the original version did not support Persian, so we modified its source code to handle Persian datasets. Our updated version is publicly available, providing the Persian community with a tool to simplify prompt creation  [^4] https://github.com/Hojjat-Mokhtarabadi/promptsource. Appendix LABEL:sec:appendix_prompts depicts an illustration of the PromptSource interface with an example of a Textual Entailment dataset. Moreover, since this system was originally integrated with Huggingface Datasets library (lhoest-etal-2021-datasets), we gathered datasets from various sources and consolidated them into a unified public repository on HuggingFace. Appendix LABEL:sec:appendix_prompts provides a sample of the crafted prompt templates for different datasets.
In addition to manual templating, we have decided to translate a subset of three question-answering datasets from the P3 collection sanh2022multitask. This decision was made to enhance the comprehensiveness and utility of our work by providing a broader scope of data. To ensure a high-quality translation, we utilized the No Language Left Behind (NLLB) (costa2022no) machine translation model, capable of single-sentence translations between 200 languages and dialects in various scripts. We employed the largest NLLB model with 3.3B parameters to achieve the best performance. A complete list of manually templated and translated datasets is given in Figure 2.
@@ 3.2 Task Augmentation and Quality Control
Instruction-tuned models are widely recognized for their significant benefits from exposure to extensive and varied tasks. Given this context, diverse prompts encompassing a range of instructions—spanning from basic to advanced language levels—are essential. Building on best practices outlined in the FLAN Collection (longpre2023flan), T0 (sanh2022multitask), and MetaICL (min-etal-2022-metaicl), we enhance task diversity by mixing and swapping different data fields within a given dataset. For example, while a dataset may initially assess a model’s ability to answer question X based on input Y, we train the model to generate question X when provided with answer Y. This approach effectively broadens the range of prompts available within a limited data pool.
@@
Table 1: List of task types, along with the number of categorization and generation prompts dedicated to each task type. The expanded version of this table can be found in the Appendix LABEL:sec:appendix_dataset_assets.

### v4

# Empowering Persian LLMs for Instruction Following:

@@ Abstract
Instruction-tuned large language models have demonstrated remarkable capabilities in following human instructions across various domains. However, their proficiency remains notably deficient in many low-resource languages. To address this challenge, we begin by introducing FarsInstruct: a comprehensive instruction dataset designed to enhance the instruction-following ability of large language models specifically for the Persian language—a significant yet underrepresented language globally. FarsInstruct encompasses a wide range of task types and datasets, each containing a mix of straightforward to complex manual written instructions, as well as translations from the Public Pool of Prompts, ensuring a rich linguistic and cultural representation. Furthermore, we introduce Co-CoLA, a framework designed to enhance the multi-task adaptability of LoRA-tuned models. Through extensive experimental analyses, our study showcases the effectiveness of the FarsInstruct dataset coupled with training by the Co-CoLA framework, in improving the performance of large language models within the Persian context. As of the current writing, FarsInstruct comprises 197 templates across 21 distinct datasets, and we intend to update it consistently, thus augmenting its applicability.
@@
Empowering Persian LLMs for Instruction Following:
A Novel Dataset and Training Approachthanks: Accepted at LoResLM @ COLING 2025
Hojjat Mokhtarabadi, Ziba Zamani, Abbas Maazallahi, Mohammad Hossein Manshaei§ Department of Electrical and Computer Engineering, Isfahan University of Technology, Iran Department of Computer Engineering, Shahid Bahonar University of Kerman, Iran Department of Electrical and Computer Engineering, University of Tehran, Tehran, Iran §Department of Computer Science, Hunter College, City University of New York, NY, USA h.mokhtarabadi@ec.iut.ac.ir, zibazamani@eng.uk.ac.ir, a.maazallahi@ut.ac.ir, mohammad.manshaei@hunter.cuny.edu
@@ 1 Introduction
The modern era of artificial intelligence is marked by numerous breakthroughs, among which is the rise of large language models (LLMs), such as GPT4 (OpenAI et al., 2024), Llama3 (Dubey et al., 2024) and PaLM (Chowdhery et al., 2022). Instruction-tuning emerges as a vital technique in the evolution of language models, involving training a model on a wide range of tasks described through natural language instructions. This method diverges from traditional task-specific fine-tuning and adapts the model’s behavior to respond to user queries with relevant and helpful answers. This technique offers a more generalized and versatile approach to model training, thus contributing significantly to the advancement of LLMs.
@@
Despite the steady progress of instruction-tuned language models, a persistent limitation remains: their difficulty in capturing the nuanced complexities of low-resource languages. This critical challenge stems from the significant gap in the availability of high-quality instruction datasets tailored to these languages. Wang et al. (2023b) highlights this concern, demonstrating that datasets lacking sufficient multilingual diversity can cause models to lose previously learned multilingual capabilities, leading to performance degradation. Moreover, translating English-centric datasets offers only partial solutions due to several inherent limitations (Naous et al., 2024; Ramesh et al., 2023; Vanmassenhove et al., 2021). While efforts have been made to compile extensive multilingual instruction datasets (Wang et al., 2022b; Singh et al., 2024; Muennighoff et al., 2022), gaps remain in creating diverse and complex prompts for languages like Persian compared to other languages.
@@
In this study, we propose FarsInstruct, a comprehensive human-annotated instruction dataset created from existing Persian NLP datasets. It includes a mixture of manually written instructions ranging from basic to proficient language levels, alongside translations from the Public Pool of Prompts (P3) (Sanh et al., 2022), which is a collection of prompted English datasets. To ensure the diversity and representativeness of FarsInstruct, we developed 197 prompt templates derived from 21 distinct public datasets. Each prompt template comprises an input template and a target template, both of which function to extract relevant data fields from their respective datasets and reformat them into a unified structure designed for the instruction-tuning objective. For example, in the case of a Textual Entailment dataset containing the fields Premise, Hypothesis, and Label, an input template might be: "Can the hypothesis be concluded from the premise? Premise: {Premise}, Hypothesis: {Hypothesis}", while a corresponding target template could be "The answer is: {label}".
@@
The collected public datasets encompass ten different task categories: Text Summarization, Textual Entailment, Text Classification, Sentiment Analysis, Word Sense Disambiguation, Query Paraphrasing, Question Answering, Reading Comprehension, Named Entity Recognition (NER), and Translation. Figure 1 depicts an instance of a prompt within our dataset after applying its respective template. A detailed overview of the FarsInsturct dataset is provided in Section 3.
@@
Figure 1: An example of the prompts utilized in the training process. The Persian version of the prompt is employed for training purposes, while the translated English version is provided to enhance comprehension. The instruction component is highlighted in black, the data fields are marked in orange, and the target answer is indicated in gray. In Appendix D, this example is shown in the PromptSource environment.
@@
Additionally, parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) (Hu et al., 2021), not only face challenges in multi-task settings but are also prone to catastrophic forgetting (Wang et al., 2023a; Li et al., 2024; Kalajdzievski, 2024). To address these issues, we propose Co-CoLA, a novel integration of CoLA (Xia et al., 2024) with rehearsal training (Kirkpatrick et al., 2017). More specifically, we adopt an iterative optimization framework that merges learned low-rank matrices into the model parameters and reinitializes optimization for new LoRA modules. At each iteration, we retrain a subset of data from previously learned tasks, mixing it with the current task’s data during training. This periodic revisiting of earlier tasks ensures that the model retains performance across both old and new tasks, all while preserving computational efficiency. Section 4 presents an in-depth explanation of the Co-CoLA method.
@@
In summary, our contributions to advancing Persian instruction understanding are threefold: (1) We present FarsInstruct, a comprehensive human-annotated instruction dataset for Persian, covering varied and representative tasks for different categories such as text summarization, named entity recognition, and translation. (2) We introduce Co-CoLA, a method that combines CoLA with rehearsal training to mitigate catastrophic forgetting in multi-task learning. (3) We release FarsInstruct as an open-source resource, with a commitment to its continued expansion to include a broader range of tasks and modalities [^1] https://huggingface.co/datasets/PNLPhub/FarsInstruct, [^2] https://github.com/Hojjat-Mokhtarabadi/FarsInstruct.
@@ 2 Related work
Instruction-tuning: Instruction tuning refers to the process of training language models using specific input-output pairs derived from diverse data sources. This approach enhances the ability of a pre-trained LLM to interpret and respond to a wide range of human requests expressed in natural language. Instruction datasets used for this purpose are typically created in one of three ways: (1) manually created by researchers from existing NLP datasets (Wang et al., 2022b; Wei et al., 2021), (2) synthesized by prompting proprietary models with a small, seed dataset (Taori et al., 2023; Wang et al., 2022a; Honovich et al., 2023), or (3) generated entirely from scratch, involving human-written prompt-response pairs (Conover et al., 2023; Köpf et al., 2024). In this work, we adopt the first approach to develop FarsInstruct. Previous works such as FLAN (Wei et al., 2021) and P3 (Sanh et al., 2022) have been instrumental in advancing instruction dataset creation. FLAN encompasses over 60 NLP datasets, while P3 features more than 2,000 prompts from 177 datasets, each significantly contributing to the field. SuperNaturalInstruction (Wang et al., 2022b) further advanced the field by assembling a comprehensive benchmark featuring 1,616 expert-written NLP tasks, covering 76 unique task types, and extending support to multiple languages. xP3 (Muennighoff et al., 2022) expanded on P3’s groundwork by including content from 46 languages, adding new tasks like Translation and Program Synthesis that P3 had not tackled. Similarly, Aya (Singh et al., 2024) represents a major multilingual effort, featuring an extensive dataset of 513 million instances across 114 languages. This was achieved through a global collaboration involving fluent speakers who contributed instructional content. Our dataset distinguishes itself from these collections in its depth and adaptability, especially with the inclusion of more challenging tasks in Persian, offering a high level of detail not found in many multilingual efforts. While most such projects primarily use machine translations and cover a narrow range of tasks, our dataset presents a wide array of culturally and linguistically rich tasks.
@@
Parameter effecient fine-tuning: Conventional full-parameter fine-tuning becomes computationally impractical as model size and the number of downstream tasks increase. To address this challenge, recent advancements in PEFT methods advocate for training only a small subset of parameters while leaving the majority of pre-trained model parameters intact. One of the most widely utilized paradigms in PEFT is Low-Rank Adaptation (LoRA) (Hu et al., 2021). LoRA modifies only a small, low-rank portion of the model’s weights by incorporating low-rank matrices into the model’s weights during the training process. Despite the significant computational advantage of LoRA, it falls short in multi-task adaptation, Additionally, Kalajdzievski (2024) demonstrated that PEFT techniques, including LoRA, remain vulnerable to catastrophic forgetting, where models lose previously acquired knowledge when fine-tuned on new tasks. MultiLoRA (Wang et al., 2023a) addresses the limitations of LoRA by reducing the dominance of top singular vectors, horizontally scaling LoRA modules, and altering the initialization of adaptation matrices, which leads to improved performance across multiple tasks with minimal additional parameters. MixLoRA (Li et al., 2024) introduces multiple LoRA-based experts within a frozen pre-trained model using a top-k routing strategy to efficiently distribute tasks, independently configure attention layer adapters, and apply auxiliary load balance loss, significantly enhancing performance while reducing GPU memory consumption and training latency. Further, CoLA (Xia et al., 2024) introduces an iterative optimization framework designed to improve the fine-tuning of LLMs by employing multiple iterations of LoRA. In this paper, we design Co-CoLA to address the issue of catastrophic forgetting, while ensuring an effective multi-task adaption.
@@ 3 FarsInstruct Dataset
With about 130 million [^3] https://en.wikipedia.org/wiki/Persian_language speakers, Persian — also referred to as Farsi in Iran — is an important language in the Middle East and Central Asia. FarsInstruct represents a project to provide a comprehensive public instruction dataset for the Persian community. As of this writing, FarsInstruct has 197 carefully designed and created prompt templates for 21 already-published public datasets and some translations from existing prompted datasets. Unlike multilingual collections focusing on common tasks such as Text Summarization and Question Answering, FarsInstruct introduces more task types, including Named Entity Recognition and Word Sense Disambiguation. The creation procedure, statistics, task augmentation, and quality of the dataset are covered in detail in the following subsections. Additional illustrations and tables are provided in the Appendix BCD.
@@ 3.1 Dataset Construction
To efficiently create a large collection of prompts, we primarily utilized PromptSource (Bach et al., 2022), an open-source tool designed for creating, sharing, and managing prompts for NLP tasks. A key design choice in Bach et al. (2022) is the use of Jinja2444https://jinja.palletsprojects.com/en/3.1.x/ as a templating language, providing the flexibility crucial for crafting clear and effective prompts. Each dataset has multiple prompt templates, each of which consists of an input and a target template. These templates map raw data fields into natural language, structuring both the input and target sequences. Practically, templates allow users to mix arbitrary text with data fields. We refer to the text within the input template that guides the model’s behavior as "Instruction". Additionally, each prompt template documents essential metadata, including evaluation metrics and the language used.
The PromptSource toolkit offers an interface for interactively writing prompts on datasets. However, the original version did not support Persian, so we modified its source code to handle Persian datasets. Our updated version is publicly available, providing the Persian community with a tool to simply create and develop prompts [^5] https://github.com/Hojjat-Mokhtarabadi/promptsource. Appendix D depicts an illustration of the PromptSource interface with an example of a Textual Entailment dataset. Moreover, since this system was originally integrated with Huggingface Datasets library (Lhoest et al., 2021), we gathered datasets from various sources and consolidated them into a unified public repository on HuggingFace. Appendix D provides a sample of the crafted prompt templates for different datasets.
In addition to manual templating, we have decided to translate a subset of three question-answering datasets from the P3 collection Sanh et al. (2022). This decision was made to enhance the comprehensiveness and utility of our work by providing a broader scope of data. To ensure a high-quality translation, we utilized the No Language Left Behind (NLLB) (Costa-jussà et al., 2022) machine translation model, capable of single-sentence translations between 200 languages and dialects in various scripts. We employed the largest NLLB model with 3.3B parameters to achieve the best performance. A complete list of manually templated and translated datasets is given in Figure 2.
@@ 3.2 Task Augmentation and Quality Control
Instruction-tuned language models are known for their significant benefits from exposure to a broad array of tasks. In this regard, we aimed to diversify the tasks through two approaches. First, we phrased the instructions at varying language levels, ranging from basic to advanced. Second, building on best practices outlined in the FLAN Collection (Longpre et al., 2023), T0 (Sanh et al., 2022), and MetaICL (Min et al., 2022), we enhanced task diversity by mixing and swapping different data fields within a given dataset. For instance, while a dataset may initially assess a model’s ability to answer question X based on input Y, we train the model to generate question X when provided with answer Y, thereby effectively broadening the range of prompts available within a limited data pool.
@@
Table 1: List of task types, along with the number of categorization and generation prompts dedicated to each task type. The expanded version of this table can be found in the Appendix C.